# TESSERA: Temporal Embeddings of Surface Spectra for Earth Representation and Analysis

### v4

@@ Acknowledgments
The pipeline for applying TESSERA embeddings to downstream tasks is as follows (illustrated in the Supplementary fig. S5):
@@
    Download Embeddings for the Region of Interest: The Geotessera Python library  (?) allows users to download embeddings for a desired region and year in the form of a numpy array.
@@
Downstream Task: Austrian Crop Classification
@@
We used the publicly available INVEKOS dataset for Austria, focusing on the 2022 growing season  (?). The data set originally contained 154 crop types, which we grouped into 17 broader classes (for example, combining sugar beet and feed beet) based on phenological similarity and sample availability to ensure robust training and evaluation. For each pixel in the dataset, we extracted its corresponding 128-dimensional TESSERA embedding from our generated 2022 global embeddings map.
@@
  • PRESTO + MLP: For a direct and fair comparison with a leading foundation model, we used the official pre-trained PRESTO model  (?) to generate its pixel embeddings. These embeddings were then fed into an MLP head identical to the one used for TESSERA.
  • GSE + MLP: To include another key foundation model, we used the publicly available Google Satellite Embedding (GSE) V1 dataset  (?). These embeddings were similarly fed into an MLP head identical to the one used for TESSERA.
@@
Patch-wise Semantic Segmentation
  • For the foundation models capable of generating pixel-wise embeddings, namely TESSERA, PRESTO  (?), and GSE  (?), a distinct workflow was used. We first construct patches from their pixel-level embeddings (e.g., a data cube of size $H\times W\times C$). These patches were then fed into a standard UNet architecture  (?), which learns to model the spatial relationships between the embeddings to produce a segmentation map. It is noteworthy that while TESSERA and PRESTO’s base embeddings do not explicitly model spatial context, GSE’s do. However, their shared ability to generate pixel-level outputs makes them suitable for this UNet-based downstream approach.
  • For other foundation models that are inherently patch-based (e.g., Prithvi  (?), Satlas  (?)), their encoders already process image patches. Therefore, for these models, we only attach and train a UPerNet decoder head to their encoders to generate the final segmentation output. We did not freeze the encoders during the fine-tuning.
Performance was measured using mean Intersection over Union (mIoU) and macro F1 scores. A visual comparison for example patches is provided in Supplementary Fig. S6.
Embedding Space Analysis
The 2D visualizations shown in the main text (Figure 3 d) were generated by applying the Uniform Manifold Approximation and Projection (UMAP)  (?) algorithm to the embeddings of every pixel within the Austrian study area. UMAP is a non-linear dimensionality reduction technique based on Riemannian geometry and algebraic topology. It constructs a weighted k-nearest neighbour graph in the original high-dimensional space, then optimizes a low-dimensional projection to primarily preserve local distances and local structure. UMAP, like other non-linear techniques such as t-SNE and LargeVis, is designed to preserve local relationships within high-dimensional data; however, it is also able to retain more of the global structure, making it particularly effective for understanding both fine-scale and broader patterns in the embedding space. In this work, the entire embedding map (for example, a shape array $H\times W\times 128$, where H and W are the height and width of the region) was used as a direct input to UMAP for the TESSERA, PRESTO, and GSE embeddings.
    $s(i)=\frac{b(i)-a(i)}{\max(a(i),b(i))}$ (S13)
    $\text{DBI}=\frac{1}{k}\sum_{i=1}^{k}\max_{j\neq i}\left(\frac{\sigma_{i}+\sigma_{j}}{d(c_{i},c_{j})}\right)$ (S14)
Downstream Task: Canopy height estimation
Study Areas and Ground Truth Datasets
[…]
Modelling methods
To predict the canopy height from the different embeddings, we used a standard U-Net architecture  (?). It takes as input a 32 $\times$ 32 pixel patch (covering a 320 $\times$ 320 m area) of representations. For 128-dimensional embeddings (as in TESSERA and PRESTO), this results in a 32 $\times$ 32 $\times$ 128 data cube. The network outputs a single-band 32 $\times$ 32 raster of predicted canopy height values. The proposed U-Net contains approximately 30 million parameters and was trained for 200 epochs using a batch size of four patches. At the end of each epoch, the model was evaluated using validation data and the model with the best performance on the validation set was retained for the final evaluation. In practice, training usually converged within the first 20 epochs, with rare boosts in performance later in training. Model training was done in Python using PyTorch.
Baselines
TESSERA, PRESTO  (?), and GSE  (?) embeddings were compared by training identical models using each representation as input, with all other settings held constant. Each model was trained for 200 epochs on the same data split, and the version achieving the best validation performance was selected for final evaluation on the held-out test set.
To benchmark performance against classical remote sensing approaches applied at the global scale, we compared the predictions against three global canopy height products: Potapov et al. [2021] (?) (GLAD), Lang et al. [2023] (?) (ETH), and Tolan et al. [2024] (?) (Meta) as well as a regional map produced by Lang et al. [2021] (?). These maps were downloaded at 10 m resolution from Google Earth Engine and compared directly to the canopy height derived from ALS. Although there are some temporal mismatches between ALS acquisition (2020) and global products, we assume relative stability of canopy height in the short term in this region, which is reasonable given that the area was not logged or burned.
Downstream Task: Burned area detection
Study Areas and Ground Truth Datasets
For fire disturbance detection, two $\times15$ km tiles in California were selected at random from the training sample pool in the burn scars benchmark dataset  (?), referred to as Burn Area A and Burn Area B. Each Burn Area contains over two million 1 m pixels. Burn Area A is a heterogeneous landscape of forests, shrublands, agricultural land, and built infrastructure. In 2020, it experienced two separate wildfires: the ’Quail Fire’ in June ($\sim$730 ha) and the much larger ’Hennessey Fire’ in August ($\sim$ 121,000 ha). Burn Area B experienced a single large fire (’Glass’) ($\sim$27,500 ha) in September 2020 that mainly affected forested land. In all cases, most of the impact of the fire occurred in vegetated areas composed of forest and shrubland. To classify the dominant land cover types in the burn areas, we used the Dynamic World Land Cover data set  (?).
The dataset from the Monitoring Trends in Burn Severity (MTBS) program  (?) was used to identify burned areas and severity of the burn. Fire ignition and containment dates, as well as fire perimeters, were obtained from the MTBS burned area boundaries dataset. Burn severity, defined as the measure of ecological damage caused by fire  (?), was derived from categorical and continuous severity metrics provided by MTBS. To obtain burn severity categories of low, medium, and high, as well as continuous burn severity, we used the annualized MTBS burn severity dataset as well as individual severity maps for each fire. MTBS calculates severity using two primary indices: the differenced NBR (dNBR) and its relativized form (RdNBR), both calculated from pre- and post-fire Landsat images. Low severity typically indicates limited ecological impact, while high severity reflects significant ecological disruption and vegetation mortality  (?). Since MTBS datasets are derived from Landsat, they are provided with a resolution of 30 m and were upsampled to 10 m to match the resolution of the TESSERA representations. The MTBS and Dynamic World datasets were downloaded as 10 m rasters from Google Earth Engine  (?). For our classification analysis, we define “unburned” pixels using MTBS class 0 and “burned” pixels as MTBS classes 2 (low), 3 (moderate) and 4 (high). The ambiguous “unburned/low” class (1) was excluded to ensure a clear binary classification benchmark.
Dimensionality Reduction
We used UMAP (Uniform Manifold Approximation and Projection)  (?) to reduce 128 dimensions of TESSERA embeddings to 2D for visual inspection of FM representations in areas affected by wildfires. We applied the dimensionality reduction to 128-dimensional TESSERA embeddings and visualized the resulting 2D projections. To interpret the structure of the projected clusters, we colour coded points according to different annotation schemes: burn state (burned vs. unburned), land cover class, or burn severity. These visualizations helped us assess whether FM embeddings spatially separate ecological conditions of interest. To reduce computational cost and improve legibility, we randomly subsampled a subset of pixels prior to applying the dimensionality reduction. UMAP projections were additionally used to classify pixels into unburned, low, moderate, and high-severity classes using a random forest (RF) classifier, to quantify the separability in the UMAP embedding space. UMAP projections were computed using the umap-learn Python library and RF was implemented using the scikit-learn Python library.
Label Efficiency Analysis and Baselines
We compared three types of input embeddings for the classification task. The first was TESSERA, using embeddings for 2020. The second was a task-specific composite baseline, constructed using prior knowledge of the fire event. This baseline consisted of median composites of 12 Sentinel-2 bands, with the Normalized Burn Ratio (NBR), and Normalized Difference Vegetation Index (NDVI) calculated separately for a pre-fire window (October 2019–April 2020) and a post-fire window (October 2020–April 2021), then concatenated into a single feature vector. The third type of input was Google Satellite Embeddings (GSE), using publicly available annual embeddings from the GSE V1 dataset  (?). GSE and task-specific composite images were downloaded as rasters from Google Earth Engine.
Downstream Task: Above-ground biomass estimation
Dataset
For the AGB regression task, we used the ground truth data from the BioMassters dataset  (?). This dataset covers nearly 11,500 forest patches in Finland, with each patch representing a $\times2,560$  m area at 10 meter pixel resolution. Above-Ground Biomass (AGB) values were determined using open forest airborne LiDAR data and an extensive network of field plots as reference measurements by the Finnish Forest Centre. Patch coordinates and collection years–provided by the authors of Ref.  (?) upon request–were used to retrieve satellite imagery to generate TESSERA and Google Satellite Embeddings (GSE)  (?).
UNet model architecture
To obtain AGB estimates, we trained a convolutional neural network based on the UNet architecture  (?), using TESSERA or GSE embeddings   (?) as input. The model received $\times256$$\times$128 and $\times256$$\times$64 input patches for TESSERA and GSE, respectively, where the dimension denotes the number of channels.
Following the evaluation protocol in Ref.  (?), RMSE was computed on a test set of 2773 patches and reported as the average RMSE per patch. For limited label testing, a subset of labels was randomly chosen from the remaining data and randomly divided into training and validation sets (80:20). The UNet model was trained on the embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model. Figure  5 shows the average RMSE calculated over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot. Training was performed with Adam Optimizer (learning rate = 1e-4, batch size = 4). Task-specific training on the entire dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours for TESSERA embeddings and 4 hours for the GSE embeddings.
Baselines
In addition to GSE, we compared the performance of TESSERA to two other baselines: the winner of the BioMassters competition  (?) and the SpectralGPT foundation model  (?) combined with a UPerNet  (?).
For SpectralGPT, we used the implementation provided in Ref.  (?), with the BioMassters satellite data as input, a lightweight spatial-temporal encoder L-TAE  (?) for temporal aggregation, and a UPerNet as the regression head. This setup was found in  (?) to outperform other tested foundation models on the BioMassters dataset. To ensure consistency in the limited label setting, we replaced the default sampling strategy with the same random splits used to evaluate TESSERA and normalized the inputs with their mean and standard deviation. Although we initially experimented with using per patch RMSE as the loss function to align with our evaluation metric, we ultimately reverted to the MSE loss  (?) because it resulted in slightly better performance. Obtaining the results for the full dataset using a single NVIDIA H200 GPU took approximately 18 hours.
Downstream Task: Stocking indices in Voluntary Carbon Markets
Relative height modelling
To model a suitable stocking index from the 128-dimensional TESSERA embeddings, we extracted 110,000 relative heights RH95 and RH10 from the GEDI Level 4A product  (?) from 2022 from the land area within the 22MCT Sentinel-2 tile in Para Sate, Brazil. From the 10 m pixel falling at the centre of each full-waveform footprint, the 128 TESSERA embeddings were extracted as predictor variables for regression analysis, with the relative height differences (dRH = RH95-RH10) used as the target variables. A random forest (RF) regression model with 500 trees and 42 variables per split was used to model this stocking index. The TerraClass Amazonia 2022 land cover map  (?) was used to distinguish between forest and non-forest. To account for implausible GEDI recordings often found in areas without significant tree cover (e.g., water surfaces or agricultural fields), the dRH values of these classes were set to zero before model training.
Stocking index baselines
For inference and assessment, the trained RF model was applied to eight Sentinel-2 tiles in the same region of Para State (MGRS tiles 22MCT, 22MCU, 22MDT, 22MDU, 22MGB, 22MGC, 22MHB, 22MHC). Within this area, a total of 38 agroforestry plots are located for which in situ measurements are available under a DTA. For comparison, five recent and widely used canopy height and AGB products were chosen and compared against the same set of 38 in situ AGB measurements: ETH CHM (10 m)  (?), CTrees canopy tree height (5 m)  (?), ESA-CCI AGB (100 m)  (?), Meta CHM (1 m)   (?), and GTDX AGB (25 m)  (?). Using the same TerraClass Amazonia 2022 shapefiles that define the agroforestry systems, the pixels of the respective products were extracted and averaged for the chronosequence plots (n=38).
  1. 1.
  2. 2.
[…]
  3. 3.
    CTrees canopy height converted using 1.806 $\times$ + 44.9
  4. 4.
    ESA AGB converted using 0.407 $\times$ + 34.0
  5. 5.
    Meta canopy height converted using 3.723 $\times$ + 46.6
  6. 6.
    GEDI-Tandem-X AGB prediction converted using 0.3 $\times$ + 54.3
Interactive Habitat Mapping Tool
Here we demonstrate our interactive classification tool, a web-based utility designed for rapid human-in-the-loop habitat mapping using pre-computed model embeddings, as seen in Figure S7. The tool allows a user to first define a region of interest, for which it fetches and mosaics the relevant high-dimensional embedding tiles. To aid visual interpretation, it performs Principal Component Analysis on the embeddings and displays the first three components as a false-colour RGB image overlay on a satellite basemap.
The core of the tool is the interactive labelling workflow. A user can define custom habitat classes and provide training data by clicking on the map to place labelled points. These points and their corresponding 128-dimensional embedding vectors are used to train a k-Nearest Neighbors (kNN) classifier. The trained model then predicts a class for every pixel in the region of interest, producing a map layer that overlays a coregistered aerial view. This process allows for iterative refinement, where a user can add more points to correct misclassifications and retrain the model for immediate feedback, demonstrating an accessible method for quickly creating habitat maps without requiring deep machine learning expertise. We also provide a batch processing tool for larger regions of interest. The tools and their open source code are freely available in our github repository  (?).
Figure S1: TESSERA is trained on globally-distributed training data. TESSERA was trained on over 3000 MGRS tiles distributed globally from 2017 to 2024. The colour of each hexagon in the map corresponds to the average number of valid observation days for Sentinel-1 (S1) and Sentinel-2 (S2), as defined by the bivariate colour legend. This visualization highlights the density of combined S1 and S2 observations available for training across different regions, ensuring the model learns from a diverse range of geographical and environmental conditions.
Figure S2: Training dynamics curves show that TESSERA rapidly reaches its asymptotes. a, Training loss curves, showing the main Barlow Twins loss, the mixup regularization loss, and the total loss. b, The learning rate schedule, consisting of a linear warmup phase followed by a cosine decay. c, RankMe score calculated on the high-dimensional (16,384-D) output of the projector, which quickly saturates, indicating the model is effectively utilizing the large feature space for redundancy reduction. d, RankMe score of the final 128-D fused representation, showing that the embeddings used for downstream tasks become progressively richer and less redundant. e, f, Downstream performance on the Austrian crop classification validation set, tracked by accuracy and F1 score respectively during pre-training, showing a steady improvement that correlates with the training progress.
Figure S3: Data shuffling greatly improves training stability. a, Schematic of the data shuffling and loading process. D-pixels (coloured squares) from thousands of MGRS tiles are first aggregated into a global pool. A custom Rust binary  (?) performs a global shuffle on this multi-terabyte dataset before applying augmentations. The processed data is then organized into chunks and streamed by data workers to form well-shuffled, globally diverse training batches. b, Comparison of training loss curves. The top plot shows the volatile loss progression typical of a localized shuffling strategy, which is susceptible to geographic bias. The bottom plot shows the significantly smoother and more stable loss curve achieved with our global shuffling pipeline, demonstrating more effective and robust model convergence.
Figure S4: Model performance improves with increasing projector size and additional computation. a, Hyperparameter sweep for the projector network’s architecture. Each point represents a configuration, with its position determined by the number of hidden layers (depth, x-axis) and their width (y-axis). The colour and size of the point indicate the validation F1 score and the number of parameters, respectively. Performance generally increases with both width and depth. The dagger (†) marks the configuration with the highest F1 score, which required a reduced batch size due to memory constraints. The star (*) indicates the final configuration selected for TESSERA v1.0, representing a balance between performance and efficiency. b, Scaling laws for TESSERA. The plot shows the validation F1 score as a function of training FLOPs and model size (in billions of parameters). A clear positive correlation is visible, indicating that performance improves as more computational resources and larger models are used. The final TESSERA v1.0 model is marked with a star (*) to demonstrate its position within this scaling trend.
Figure S5: Unlike common Remote Sensing Foundation Models (RSFM), TESSERA simplifies downstream task applications. Most existing RSFMs use approach (a) where, after training, the encoder is extracted and combined with a task-specific head to form a new model for downstream tasks. Users can then fine-tune the entire model or only the head. Another approach (b) is where both the encoder and the projector (or decoder) are typically updated. In contrast, the TESSERA approach (c) uses pre-generated embeddings directly as input. This results in lightweight downstream models that consist only of a task-specific head.
Figure S6: TESSERA predictions most closely resemble the ground truth for patch-based semantic segmentation. Each row corresponds to a different patch size (32x32, 64x64, 128x128). The first column shows the ground truth segmentation. For each model (TESSERA, Presto, and GSE), two columns are displayed: a PCA visualization of the input embeddings for the patch (labelled ’Embedding’) and the final predicted segmentation map (labelled ’Predicted’).
Figure S7: We provide an interactive tool to allow non-specialist users to map TESSERA embeddings to user-defined classes. We show a PCA-derived false colour map overlaid on an RGB aerial image of South Cambridge.
Table S1: Variable definitions for remote sensing data structure
Variable Dimensions Description
$R_{t}$ $W\times H\times C$ Remote sensing data tile at time $t$
$\mathbf{D}$ $T\times W\times H\times C$ Complete temporal data stack
$\mathbf{M}$ $T\times W\times H$ Complete temporal mask stack
$S_{i,j,c}$ $T\times 1$ Time series for channel $c$ at location $(i,j)$
$W,H$ Scalar Spatial dimensions (width, height)
$C$ Scalar Number of spectral channels + polarizations
$T$ Scalar Number of time steps
Table S2: Architectural parameters of the TESSERA v1.0 model.
Component Parameters Percentage of Total
Encoder (Feature Extractor) 45,697,026 3.29%
   Sentinel-1 Encoder 22,846,465 1.64%
   Sentinel-2 Encoder 22,850,561 1.64%
Fusion MLP 131,200 <<0.01%
Projector (for Training only) 1,344,700,421 96.71%
Total Parameters 1,390,528,647 100.00%
Table S3: Pre-training hyperparameters for the TESSERA v1.0 model.
@@
Table S4: The ablation study demonstrates that data shuffling and mixup regularization are essential, Sentinel 1 data adds little additional value, and that quantization has nearly no impact on performance. We study downstream crop classification performance on the 100x-downsampled Austrian crop dataset. The baseline represents the full model. Each subsequent row shows the result of removing the specified component. Performance is measured by validation F1 score and the RankMe score. Arrows ($\uparrow$) indicate that higher values are better. The results demonstrate that all components contribute positively to the model’s overall performance.
Baseline (Full Model) 0.773 0.963
w/o Global Data Shuffling 0.681 0.847
w/o Mixup Regularization 0.662 0.857
w/o Sentinel-1 Data 0.742 0.931
w/o Shuffling & Mixup 0.626 0.867
w/o Embedding Quantization 0.779 0.972

### v5

@@ A.3 Downstream Task Application Methodology
The pipeline for applying TESSERA embeddings to downstream tasks is as follows (illustrated in the Supplementary fig. 12):
@@
    Download Embeddings for the Region of Interest: The Geotessera Python library [37] allows users to download embeddings for a desired region and year in the form of a numpy array.
@@
A.4 Downstream Task: Austrian Crop Classification
@@
We used the publicly available INVEKOS dataset for Austria, focusing on the 2022 growing season [44]. The data set originally contained 154 crop types, which we grouped into 17 broader classes (for example, combining sugar beet and feed beet) based on phenological similarity and sample availability to ensure robust training and evaluation. For each pixel in the dataset, we extracted its corresponding 128-dimensional TESSERA embedding from our generated 2022 global embeddings map.
@@
  • PRESTO + MLP: For a direct and fair comparison with a leading foundation model, we used the official pre-trained PRESTO model [19] to generate its pixel embeddings. These embeddings were then fed into an MLP head identical to the one used for TESSERA.
  • GSE + MLP: To include another key foundation model, we used the publicly available Google AphaEarth Satellite Embedding (GSE) V1 dataset [39]. These embeddings were similarly fed into an MLP head identical to the one used for TESSERA.
@@
Table 4: Pre-training hyperparameters for the TESSERA v1.0 model.
@@
Patch-wise Semantic Segmentation
  • For the foundation models capable of generating pixel-wise embeddings, namely TESSERA, PRESTO [19], and GSE [39], a distinct workflow was used. We first construct patches from their pixel-level embeddings (e.g., a data cube of size $H\times W\times C$). These patches were then fed into a standard UNet architecture [76], which learns to model the spatial relationships between the embeddings to produce a segmentation map. It is noteworthy that while TESSERA and PRESTO’s base embeddings do not explicitly model spatial context, GSE’s do. However, their shared ability to generate pixel-level outputs makes them suitable for this UNet-based downstream approach.
  • For other foundation models that are inherently patch-based (e.g., Prithvi [110], Satlas [15]), their encoders already process image patches. Therefore, for these models, we only attach and train a UPerNet decoder head to their encoders to generate the final segmentation output. We did not freeze the encoders during the fine-tuning.
Performance was measured using mean Intersection over Union (mIoU) and macro F1 scores. A visual comparison for example patches is provided in Supplementary Fig. 13.
Figure 13: TESSERA predictions most closely resemble the ground truth for patch-based semantic segmentation. Each row corresponds to a different patch size (32x32, 64x64, 128x128). The first column shows the ground truth segmentation. For each model (TESSERA, Presto, and GSE), two columns are displayed: a PCA visualization of the input embeddings for the patch (labelled ’Embedding’) and the final predicted segmentation map (labelled ’Predicted’).
Embedding Space Analysis
The 2D visualizations shown in the main text (Figure 3 d) were generated by applying the Uniform Manifold Approximation and Projection (UMAP) [45] algorithm to the embeddings of every pixel within the Austrian study area. UMAP is a non-linear dimensionality reduction technique based on Riemannian geometry and algebraic topology. It constructs a weighted k-nearest neighbour graph in the original high-dimensional space, then optimizes a low-dimensional projection to primarily preserve local distances and local structure. UMAP, like other non-linear techniques such as t-SNE and LargeVis, is designed to preserve local relationships within high-dimensional data; however, it is also able to retain more of the global structure, making it particularly effective for understanding both fine-scale and broader patterns in the embedding space. In this work, the entire embedding map (for example, a shape array $H\times W\times 128$, where H and W are the height and width of the region) was used as a direct input to UMAP for the TESSERA, PRESTO, and GSE embeddings.
    $s(i)=\frac{b(i)-a(i)}{\max(a(i),b(i))}$ (13)
    $\text{DBI}=\frac{1}{k}\sum_{i=1}^{k}\max_{j\neq i}\left(\frac{\sigma_{i}+\sigma_{j}}{d(c_{i},c_{j})}\right)$ (14)
A.5 Downstream Task: Canopy height estimation
Study Areas and Ground Truth Datasets
[…]
Modelling methods
To predict the canopy height from the different embeddings, we used a standard U-Net architecture [76]. It takes as input a 32 $\times$ 32 pixel patch (covering a 320 $\times$ 320 m area) of representations. For 128-dimensional embeddings (as in TESSERA and PRESTO), this results in a 32 $\times$ 32 $\times$ 128 data cube. The network outputs a single-band 32 $\times$ 32 raster of predicted canopy height values. The proposed U-Net contains approximately 30 million parameters and was trained for 200 epochs using a batch size of four patches. At the end of each epoch, the model was evaluated using validation data and the model with the best performance on the validation set was retained for the final evaluation. In practice, training usually converged within the first 20 epochs, with rare boosts in performance later in training. Model training was done in Python using PyTorch.
Baselines
TESSERA, PRESTO [19], and GSE [39] embeddings were compared by training identical models using each representation as input, with all other settings held constant. Each model was trained for 200 epochs on the same data split, and the version achieving the best validation performance was selected for final evaluation on the held-out test set.
To benchmark performance against classical remote sensing approaches applied at the global scale, we compared the predictions against three global canopy height products: Potapov et al. [2021][59] (GLAD), Lang et al. [2023][58] (ETH), and Tolan et al. [2024][60] (Meta) as well as a regional map produced by Lang et al. [2021][68]. These maps were downloaded at 10 m resolution from Google Earth Engine and compared directly to the canopy height derived from ALS. Although there are some temporal mismatches between ALS acquisition (2020) and global products, we assume relative stability of canopy height in the short term in this region, which is reasonable given that the area was not logged or burned.
A.6 Downstream Task: Burned area detection
Study Areas and Ground Truth Datasets
For fire disturbance detection, two $\times15$ km tiles in California were selected at random from the training sample pool in the burn scars benchmark dataset [114], referred to as Burn Area A and Burn Area B. Each Burn Area contains over two million 1 m pixels. Burn Area A is a heterogeneous landscape of forests, shrublands, agricultural land, and built infrastructure. In 2020, it experienced two separate wildfires: the ’Quail Fire’ in June ($\sim$730 ha) and the much larger ’Hennessey Fire’ in August ($\sim$ 121,000 ha). Burn Area B experienced a single large fire (’Glass’) ($\sim$27,500 ha) in September 2020 that mainly affected forested land. In all cases, most of the impact of the fire occurred in vegetated areas composed of forest and shrubland. To classify the dominant land cover types in the burn areas, we used the Dynamic World Land Cover data set [115].
The dataset from the Monitoring Trends in Burn Severity (MTBS) program [95] was used to identify burned areas and severity of the burn. Fire ignition and containment dates, as well as fire perimeters, were obtained from the MTBS burned area boundaries dataset. Burn severity, defined as the measure of ecological damage caused by fire [116], was derived from categorical and continuous severity metrics provided by MTBS. To obtain burn severity categories of low, medium, and high, as well as continuous burn severity, we used the annualized MTBS burn severity dataset as well as individual severity maps for each fire. MTBS calculates severity using two primary indices: the differenced NBR (dNBR) and its relativized form (RdNBR), both calculated from pre- and post-fire Landsat images. Low severity typically indicates limited ecological impact, while high severity reflects significant ecological disruption and vegetation mortality [116]. Since MTBS datasets are derived from Landsat, they are provided with a resolution of 30 m and were upsampled to 10 m to match the resolution of the TESSERA representations. The MTBS and Dynamic World datasets were downloaded as 10 m rasters from Google Earth Engine [102]. For our classification analysis, we define “unburned” pixels using MTBS class 0 and “burned” pixels as MTBS classes 2 (low), 3 (moderate) and 4 (high). The ambiguous “unburned/low” class (1) was excluded to ensure a clear binary classification benchmark.
Dimensionality Reduction
We used UMAP (Uniform Manifold Approximation and Projection) [45] to reduce 128 dimensions of TESSERA embeddings to 2D for visual inspection of FM representations in areas affected by wildfires. We applied the dimensionality reduction to 128-dimensional TESSERA embeddings and visualized the resulting 2D projections. To interpret the structure of the projected clusters, we colour coded points according to different annotation schemes: burn state (burned vs. unburned), land cover class, or burn severity. These visualizations helped us assess whether FM embeddings spatially separate ecological conditions of interest. To reduce computational cost and improve legibility, we randomly subsampled a subset of pixels prior to applying the dimensionality reduction. UMAP projections were additionally used to classify pixels into unburned, low, moderate, and high-severity classes using a random forest (RF) classifier, to quantify the separability in the UMAP embedding space. UMAP projections were computed using the umap-learn Python library and RF was implemented using the scikit-learn Python library.
Label Efficiency Analysis and Baselines
We compared three types of input embeddings for the classification task. The first was TESSERA, using embeddings for 2020. The second was a task-specific composite baseline, constructed using prior knowledge of the fire event. This baseline consisted of median composites of 12 Sentinel-2 bands, with the Normalized Burn Ratio (NBR), and Normalized Difference Vegetation Index (NDVI) calculated separately for a pre-fire window (October 2019–April 2020) and a post-fire window (October 2020–April 2021), then concatenated into a single feature vector. The third type of input was Google AlphaEarth Satellite Embeddings (GSE), using publicly available annual embeddings from the GSE V1 dataset [39]. GSE and task-specific composite images were downloaded as rasters from Google Earth Engine.
A.7 Downstream Task: Above-ground biomass estimation
Dataset
For the AGB regression task, we used the ground truth data from the BioMassters dataset [72]. This dataset covers nearly 11,500 forest patches in Finland, with each patch representing a $\times2,560$  m area at 10 meter pixel resolution. Above-Ground Biomass (AGB) values were determined using open forest airborne LiDAR data and an extensive network of field plots as reference measurements by the Finnish Forest Centre. Patch coordinates and collection years–provided by the authors of Ref. [72] upon request–were used to retrieve satellite imagery to generate TESSERA and Google AlphaEarth Satellite Embeddings (GSE) [39].
UNet model architecture
To obtain AGB estimates, we trained a convolutional neural network based on the UNet architecture [76], using TESSERA or GSE embeddings  [39] as input. The model received $\times256$$\times$128 and $\times256$$\times$64 input patches for TESSERA and GSE, respectively, where the dimension denotes the number of channels.
Following the evaluation protocol in Ref. [72], RMSE was computed on a test set of 2773 patches and reported as the average RMSE per patch. For limited label testing, a subset of labels was randomly chosen from the remaining data and randomly divided into training and validation sets (80:20). The UNet model was trained on the embeddings for 80 epochs, with the model achieving the lowest loss on the validation set selected as the final model. Figure  5 shows the average RMSE calculated over five random splits, with bars indicating the minimum–maximum range when it exceeds the size of the dot. Training was performed with Adam Optimizer (learning rate = 1e-4, batch size = 4). Task-specific training on the entire dataset (80% training / 20% validation) using a single NVIDIA H200 GPU and 8 CPU cores took approximately 5 hours for TESSERA embeddings and 4 hours for the GSE embeddings.
Baselines
In addition to GSE, we compared the performance of TESSERA to two other baselines: the winner of the BioMassters competition [72] and the SpectralGPT foundation model [18] combined with a UPerNet [78].
For SpectralGPT, we used the implementation provided in Ref. [79], with the BioMassters satellite data as input, a lightweight spatial-temporal encoder L-TAE [77] for temporal aggregation, and a UPerNet as the regression head. This setup was found in [79] to outperform other tested foundation models on the BioMassters dataset. To ensure consistency in the limited label setting, we replaced the default sampling strategy with the same random splits used to evaluate TESSERA and normalized the inputs with their mean and standard deviation. Although we initially experimented with using per patch RMSE as the loss function to align with our evaluation metric, we ultimately reverted to the MSE loss [79] because it resulted in slightly better performance. Obtaining the results for the full dataset using a single NVIDIA H200 GPU took approximately 18 hours.
A.8 Downstream Task: Stocking indices in Voluntary Carbon Markets
Relative height modelling
To model a suitable stocking index from the 128-dimensional TESSERA embeddings, we extracted 110,000 relative heights RH95 and RH10 from the GEDI Level 4A product [53] from 2022 from the land area within the 22MCT Sentinel-2 tile in Para Sate, Brazil. From the 10 m pixel falling at the centre of each full-waveform footprint, the 128 TESSERA embeddings were extracted as predictor variables for regression analysis, with the relative height differences (dRH = RH95-RH10) used as the target variables. A random forest (RF) regression model with 500 trees and 42 variables per split was used to model this stocking index. The TerraClass Amazonia 2022 land cover map [117] was used to distinguish between forest and non-forest. To account for implausible GEDI recordings often found in areas without significant tree cover (e.g., water surfaces or agricultural fields), the dRH values of these classes were set to zero before model training.
Stocking index baselines
For inference and assessment, the trained RF model was applied to eight Sentinel-2 tiles in the same region of Para State (MGRS tiles 22MCT, 22MCU, 22MDT, 22MDU, 22MGB, 22MGC, 22MHB, 22MHC). Within this area, a total of 38 agroforestry plots are located for which in situ measurements are available under a DTA. For comparison, five recent and widely used canopy height and AGB products were chosen and compared against the same set of 38 in situ AGB measurements: ETH CHM (10 m) [58], CTrees canopy tree height (5 m) [118], ESA-CCI AGB (100 m) [119], Meta CHM (1 m)  [60], and GTDX AGB (25 m) [120]. Using the same TerraClass Amazonia 2022 shapefiles that define the agroforestry systems, the pixels of the respective products were extracted and averaged for the chronosequence plots (n=38).
  1. 1.
  2. 2.
[…]
  3. 3.
    CTrees canopy height converted using 1.806 $\times$ + 44.9
  4. 4.
    ESA AGB converted using 0.407 $\times$ + 34.0
  5. 5.
    Meta canopy height converted using 3.723 $\times$ + 46.6
  6. 6.
    GEDI-Tandem-X AGB prediction converted using 0.3 $\times$ + 54.3
A.9 Interactive Habitat Mapping Tool
Figure 14: We provide an interactive tool to allow non-specialist users to map TESSERA embeddings to user-defined classes. We show a PCA-derived false colour map overlaid on an RGB aerial image of South Cambridge.
Here we demonstrate our interactive classification tool, a web-based utility designed for rapid human-in-the-loop habitat mapping using pre-computed model embeddings, as seen in Figure 14. The tool allows a user to first define a region of interest, for which it fetches and mosaics the relevant high-dimensional embedding tiles. To aid visual interpretation, it performs Principal Component Analysis on the embeddings and displays the first three components as a false-colour RGB image overlay on a satellite basemap.
The core of the tool is the interactive labelling workflow. A user can define custom habitat classes and provide training data by clicking on the map to place labelled points. These points and their corresponding 128-dimensional embedding vectors are used to train a k-Nearest Neighbors (kNN) classifier. The trained model then predicts a class for every pixel in the region of interest, producing a map layer that overlays a coregistered aerial view. This process allows for iterative refinement, where a user can add more points to correct misclassifications and retrain the model for immediate feedback, demonstrating an accessible method for quickly creating habitat maps without requiring deep machine learning expertise. We also provide a batch processing tool for larger regions of interest. The tools and their open source code are freely available in our github repository [37].